Can you help solve these functions with 3 variables?
4a + 2b - c = 5
2a+ b - 5c = -11
a - 2b + 3c = 6

Found 3 solutions by Alan3354, mananth, Edwin McCravy:
Answer by Alan3354(69443)

Can you help solve these functions with 3 variables?
4a + 2b - c = 5
2a+ b - 5c = -11
a - 2b + 3c = 6
-----------
Maybe.
I suggest multiplying the 2nd eqn by 2, then adding or subtracting it to the 1st and 3rd eqns to eliminate b.
Then you'll have 2 eqns in a & c.

Answer by mananth(16949)

4a+-2b- c = 5 -------------- 1
2a+5b- c = -11 -------------- 2
a-2b+3c =6 -------------- 3

consider equation 1 &2 Eliminate b
Multiply 1 by 5
Multiply 2 by 2
we get
20a -10 b + -5 c = 25
4a+ 10 b + -2 c = -22
Add the two
24 a + 0 b + -7 c = 3 ------------- 4
consider equation 2 & 3 Eliminate b
Multiply 2 by 2
Multiply 3 by 5
we get
4 a + 10 b + -2 c = -22
5 a + -10 b + 15 c = 0
Add the two
9a +0b+13 c = -22 -------------5 5
Consider (4) & (5) Eliminate a
Multiply 4 by 9
Multiply (5) by -24
we get 2
216 a + -63 c = 27
-216 a + -312 c = 528
Add the two
0 a + -375 c = 555
/ -375
c = -1.48 => -37/25
Plug the value of c in (5)
9 a + -19.24 = -22
9 a = -2.76
a = -0.31 OR - 23/75
plug value of x & z in 1
-1.23 + -2 b + 1.48 = 5
-2 b = 1.23 + -1.48 + 5
b = -2.37 OR -19/8


Answer by Edwin McCravy(20086)

eq. 1       4a + 2b -  c =   5
eq. 2       2a +  b - 5c = -11
eq. 3        a - 2b + 3c =   6

You can eliminate b immediately from eqs. 1 and 3 by adding them
vertically term-by-term:

eq. 1       4a + 2b -  c =   5
eq. 3        a - 2b + 3c =   6
           -------------------
eq. 4       5a      + 2c =  11

So we must eliminate be from another pair of equations:

We multiply eq. 2 through by 2 

            4a + 2b - 10c = -22

so we can add eq. 3 to it and eliminate b from that pair:

            4a + 2b - 10c = -22
eq. 3        a - 2b +  3c =   6
            -------------------
eq. 5       5a      -  7c = -16

Now we take eqs. 4 and 5 together as a system of two equations in two
unknowns:

eq. 4       5a + 2c =  11
eq. 5       5a - 7c = -16
   
We multiply eq. 5 by -1 

           -5a + 7c = 16

so we can add eq. 4 to it to eliminate "a":

           -5a + 7c = 16
eq. 4       5a + 2c = 11
          ----------------
                 9c = 27
                  c = 3
   
Substitute c = 3 in eq. 4

          5a + 2(3) =  11
             5a + 6 = 11
                 5a = 5
                  a = 1

Substitute c = 3 and a = 1 in eq. 3

    (1) - 2b + 3(3) =   6
         1 - 2b + 9 = 6
            10 - 2b = 6
                -2b = -4
                  b = 2

So the solution is (a,b,c) = (1,2,3)

Edwin